OpenStudy (cocoacat):

Hi! Need help on a question: A 65.0 L stainless steel container was charged with 1.00 atm of hydrogen gas and 2.00 atm of oxygen gas. A spark ignited the mixture, producing water. What is the pressure in the tank at a constant temperature of 25°C? I got 1.50 atm for this part - this was correct! This is the part I need help with: What is the pressure in the tank, starting with the same pressures of gases, at a constant temperature of 125°C?

OpenStudy (diglet):

Quick question. Is the 1.5atm you calculated, the value for the pressure of the water vapour after combustion or before combustion?

OpenStudy (cocoacat):

After combustion!

OpenStudy (diglet):

okay, not 100% sure on this answer but whatever. Pressure is directly proportional to temperature (P=kT) where k is a constant of the gases above. where T is the temperature in kelvin. So first calculate k for the first part k = P/T. k= 1.5atm / (273+25) (273 + 25) is temperature in kelvin k= about 0.005 then use this constant to calculate the pressure in the second part. P = kxT P = 0.005 x (273 +125) P=1.99 atm. I used rounded values thought so be careful.

OpenStudy (cuanchi):

just use P1/T1=P2/T2 the volume is constant and the number of moles too